OpenStudy (anonymous):

Square Root of 25... 2 answers

OpenStudy (ivettef365):

what is sqrt 25 = ??

OpenStudy (anonymous):

5?

OpenStudy (ivettef365):

ok, so it could be 5 and -5

OpenStudy (anonymous):

Square roots always have a positive and negative answer since \(5\times 5= 25\) and \(-5 \times -5 = 25\)

OpenStudy (calculusxy):

What two numbers (that are the same) give you the product of 25?

OpenStudy (anonymous):

A.13 B.(-25)^1/2 C.25^1/2 D.-5

OpenStudy (anonymous):

square roots are the same as raising a number to the 1/2 power. cube roots are the same as raising a number to the 1/3 powers. -exponents means to invert the fraction.

OpenStudy (anonymous):

Which of the following are SqRt of the number below?

OpenStudy (anonymous):

\[\sqrt{x}=x^{\frac{1}{2}}\] Which do you think it is?